Can someone please help me with this question i need a worked problem as an example and an explaination....Give an example of when and why one would use a continuity correction factor?
Answer by stanbon(75887)

Here's an example someone else requested.
---
Regarding your "Why" question, please use Google
to find out what is said on that topic online.
===================================================

Find the normal approximation for the binomial probability that x = 6,
where n = 15 and p = 0.3. Compare this probability to the value of P(x=6)
----------
u = np = 15*0.3 = 4.5
s = sqrt(4.5*0.7) = 1.7748
---------------------------------------
P(x = 6) = P(5.5 < x < 6.5)
---
z(5.5) = (5.5-4.5)/1.7748 = 0.5634
z(6.5) = (6.5-4.5)/1.7748 = 1.1269
---------------
P(5.5 < x < 6.5) = P(0.5634< z <1.1269) = 0.1567
====================
Using the binomial distribution:
P(x = 5) = binompdf(15,0.3,5) = 0.2061
============================
Cheers,
Stan H.